# scenarioparampadding=8pt, margin-top=3pt, margin-bottom=3pt, border-style=none, background-color=quotebackground

### v1

@@ 3. Formative Study: Focus Group Interview
To inform the design of MindfulDiary, we first conducted a Focus Group Interview (FGI) with MHPs. We recruited six MHPs (E1–6; two males and four females)—four clinical psychologists and two psychiatrists—from a local university hospital in Korea. We held two one-hour remote sessions via Zoom. In the FGI, we first provided an overview of language model technologies and LLM’s natural language understanding and generation capabilities. We also described the current challenges of LLMs, such as potential errors and hallucinations. We then asked participants a focused set of questions on (1) the challenges MHPs currently face during patient treatment and counseling sessions and (2) their expectations and envisioned opportunities of LLMs’ role in clinical mental health settings. The session was video recorded and later transcribed. We open-coded the transcripts to identify emerging themes. In the following, we cover the findings from the FGI.
3.1. Challenges in Eliciting Responses from Patients with Depression
@@
Participants also noted that they had their patients engaged in paper-based diary writing programs but most demonstrated low participation rates and low engagement: “We tried the diary program on paper(in the inpatient ward), and several patients did write. What we saw was quite trivial, like, ‘I just felt bad today.’ But we learned there were significant events upon consultation, like having a big argument with other patients, which they did not record. Because patients with depression, or those who have had suicidal or self-harming incidents, often have a dulled state in expressing their emotions or feel apathetic, they tend to find such expressions very difficult.” (E3)
Participants envisioned that language models have significant potential given their natural and flexible conversational abilities. They suggested that this ability could help patients explore their daily lives and engage in deep reflection, making the daily logging process more engaging by providing a sense of companionship.
3.2. The Meaning of Daily Records as Perceived by MHPs
Our participants emphasized that patient records about their daily experiences can serve as crucial data for MHPs to understand them. E5 envisioned that such records could be utilized to understand a patient’s automatic thoughts and reaction mechanisms concerning specific events. For instance, in a doctor’s office, a patient with social phobia might briefly say, “I feel scared in crowded places,” and the conversation might not delve much deeper. However, the true depth of this fear can be more accurately and vividly captured if the patient records their feelings immediately after exposure to a crowded place or even later that night.
Participants imagined that language model outputs about patient journaling patterns could provide valuable insights about what counseling styles they should use for individual patients: “Observing how they respond to certain inputs can be valuable data. Some patients feel comforted by mere reassurance, while others prefer more direct, pointed feedback.”
@@ 3.2. The Meaning of Daily Records as Perceived by MHPs
From the FGI, we refined the initial concept of the MindfulDiary program. We leveraged the conversational abilities of LLMs to help patients document their daily experiences between clinical visits. MHPs had access to the collected data to inform their clinical decision-making. Furthermore, both MHPs and the research team concur that LLMs should not act solely as the primary intervention due to their inherent limitations but should function as supportive tools for clinical consultations. The subsequent section outlines the design and development process of our system.
@@ 4. MindfulDiary
{longfbox}[scenarioparam]Jane, diagnosed with chronic anxiety, frequently grapples with panic attacks. To keep track of her daily experiences, her psychiatrist recommends trying MindfulDiary as part of her treatment plan.
Every evening, Jane records her daily activities, emotions, and thoughts using the MindfulDiary app. The AI leads the conversation by asking prompt questions about the day of Jane. Additionally, the system transforms Jane’s inputs into a journal-style format, in which she can organize thoughts and reflect on them later. She can easily read the summarized content whenever she wishes to look back at past events or thoughts.
@@
4.1. MindfuleDiary App
The MindfulDiary app for patients aims to support people who might have difficulty journaling due to apathy and cognitive load through naturalistic conversation driven by an LLM. The app consists of a home screen containing an introduction and guide to the program (Figure 2), a journal writing screen (Figure 2, 2), and a screen to review the diary entries (Figure 2).
@@ 4.1.1. Journaling User Interface
Figure 3 illustrates the overall use flow of the journaling session, which begins with a Pre-Journaling Assessment (Figure 3-1⃝) that asks to fill out a questionnaire for mental health. The questionnaire comprised the modified PHQ-9 (Kroenke et al., 2001) and a custom open-ended question inquiring about recent attempts of self-harm or suicide. This assessment prevents users who provided any clues of suicidal or self-harm from journaling on the same day. (We cover this feature in detail in Section 5.4 Ethical Consideration.)
@@ 4.1.2. Conversation Design
As a result, we designed the conversation of a journaling session to follow a sequence of three stages: Rapport building, Exploration, and Wrap-up. The Rapport Building state is an ice-breaker, centered on casual exchanges about a user’s day. In this state, the assistant also shares bits of information to encourage users’ openness. As we progress to the Exploration state, the emphasis shifts to a comprehensive understanding of the user’s daily events, feelings, and thoughts, facilitated by a mix of open-ended and closed-ended queries that ensure users remain engaged and in control of the dialogue. The conversation then transitions to the Wrap-up, emphasizing completion and ensuring users have fully voiced their experiences while the system remains empathetic and receptive to any lingering topics.
Besides the three main stages, we also incorporated the Sensitive Topic state that handles the most sensitive subjects, such as self-harm or suicidal ideation. When this state is triggered, psychiatrists receive instant notifications. This allows them to oversee the conversation in real-time and step in to assist the patient if necessary. Here, the system begins by empathizing with the user, recognizing their struggles, and offering a reassuring message. Following this, the system gently probes the depth of their suicidal or self-harm thoughts. If the user expresses intense or specific plans related to self-harm or suicide, the system urges them to seek prompt assistance, either at a hospital emergency room or via the local helpline.
@@ 4.1.3. Conversational Pipeline
Lengthy and complex input prompts for LLMs are known to cause poor task performance (Brown et al., 2020) by partly omitting latent concepts (Wu et al., 2022). To steer the LLM to comply with the conversational design we intended diligently, we designed MindfulDiary’s dialogue system as a state machine. Each conversation stage is carried on with a dedicated input prompt, which is more succinct and clear than a single master prompt containing instructions for all stages. Figure 4 illustrates our conversation pipeline that runs each time a new user message is received. The pipeline incorporates two LLM-driven modules, a dialogue analyzer and the response generator. The dialogue analyzer handles the stage transition, returning the stage suggestion—whether to stay or move to a new stage—and a summary paragraph of the current dialogue from the current dialogue information. The system then formulates an LLM prompt, combining a prompt for the current stage, the dialogue summary, and the recent six messages. Receiving the prompt as an input, the response generator generates an AI message.
@@ 4.3. Clinician Dashboard
Journal. This section displays the content of the journals written by patients. The information is presented in a card format, where each card offers a summary of the journal, including timestamps, total time taken to write the journal, and associated PHQ-9 score. The interaction logs between the patient and MindfulDiary are also provided in this section.
Insights. To assist professionals in browsing through the diary, this section visualizes (1) a word cloud to understand frequent terms that the participant used at a glance, (2) a summary of major events to highlight significant happenings and (3) summary of emotions to gauge the mood based on user input. When a specific period is selected for review, a comprehensive summary is generated. This summary relies on the combined power of gpt-4 and a Korean morphological analysis tool named Kiwi (Lee, 2022). Due to the limitations of language model-driven analysis, there might be occasional inaccuracies in the generated content. First-time users of this interface are alerted about possible inaccuracies. An in-interface tooltip also reminds users that the summarized outcomes might not be accurate.
@@ 5. Field Deployment Study
Figure 5. Procedure of the Four-Week Field Deployment Study: A four-week exploration into the utilization of MindfulDiary by outpatient patients, encompassing daily use, and its integration into clinical decision-making. *Some participants did not have a follow-up visit during the experimental period
Through the four-week field deployment study with 28 patients undergoing outpatient treatment, we aimed to explore how patients and MHPs utilize MindfulDiary and what opportunities and challenges arise from its real-world use. Figure 5 illustrates the process of field deployment study.
@@ 5.1. Recruitment
We targeted outpatients from the Department of Mental Health at a University Hospital. Participants were selected based on specific criteria: (1) those who had been diagnosed with MDD and (2) those who did not exhibit heightened impulsive tendencies or harbor specific intentions towards self-harm or suicide. Eligible participants were identified through evaluations conducted by psychiatrists. Flyers and consent forms were distributed to eligible patients. For minors, the consent form process was adhered to only when they were accompanied by a guardian at the hospital.
We provided a compensation of approximately 60 USD (80,000 KRW) to participants who completed the entire four-week study process. For those who dropped out midway, we provided compensation based on the number of weeks they participated. Participants received approximately 11 USD (15,000 KRW) for every week that they were involved in (e.g., if someone participated for ten days, they would receive 11 USD). Participants were informed that consistent non-use of the system would be a criterion for study dropout. Specifically, if a participant did not engage with the MindfulDiary for three consecutive days, they would receive a notification from the researcher encouraging its use. However, continued non-use after this reminder would result in the participant being considered non-compliant with the study protocol, leading to their withdrawal from the study. This decision was made to ensure data reliability.
Through this process, a total of 36 individuals registered, of whom 8 dropped out, resulting in a final count of 28 participants who completed the 4-week field study. The majority of participants were adolescents and adults, with ages ranging from 12 to 28 years, with an average age of 17.6 ($SD=3.26$). Among the 28 participants, 11 were male, and 17 were female.
5.2. Data Collection
Patient-MindfulDiary Interaction Data. With the consent of the participants, we collected all data from their interactions with the MindfulDiary along with the raw input content and outputs from the language model.

### v2

@@ 3. Formative Study: Focus Group Interview
To inform the design of MindfulDiary, we first conducted a Focus Group Interview (FGI) with MHPs. The goal of the FGI was to understand MHPs’ perspectives, expectations, values, and precautions in utilizing LLMs in the clinical mental health context. Based on this understanding, we aimed to design the functions and interactions that the system should provide. This was an essential process in our overall approach, not just technology-centered system design, but creating a system meaningful to users and stakeholders (Thieme et al., 2022).
3.1. Procedure and Analysis
We distributed recruitment flyers in the Department of Psychiatry at a local university hospital, inviting Mental Health Professionals (MHPs) working in departments of psychiatry and mental health care centers to participate. We recruited six MHPs (E1–6; two males and four females)—four clinical psychologists and two psychiatrists whose careers varied from 1 to 11 years. Four were clinical psychologists responsible for counseling and daily monitoring and intervention of at-risk patient groups in local mental health centers and university hospitals, and two were psychiatrists in charge of outpatient and inpatient ward treatment in the psychiatry department of university hospitals (see Table 1).
We invited participants to two 1-hour remote sessions on Zoom. Two researchers participated in the sessions. We first provided an overview of language model technologies and LLM’s natural language understanding and generation capabilities until we shared a common understanding of the principles, applications, opportunities, and limitations of LLMs. Considering that we were designing a system for individuals with mental health challenges, we thoroughly covered the drawbacks of LLMs, such as uncertainty in control and hallucinations.
After the overview, we went through group discussions on how LLMs could be utilized in the current patient treatment process. As a probe, we asked participants a focused set of questions on (1) the challenges MHPs currently face during patient treatment and counseling sessions, and (2) their expectations and envisioned opportunities of LLMs’ role in clinical mental health settings. We sought to understand the experts’ perspectives through questions such as, ‘What are the difficulties or challenges patients face in their daily lives between treatments (or counseling)?’, ‘What are the important considerations in self-care that patients perform in their daily lives?’, and ‘What questions or conversational techniques do you use to encourage patients to share about their daily lives and moods?’. The session was video recorded and later transcribed. We open-coded the transcripts to identify emerging themes. In the following, we cover the findings from the FGI.
Table 1. Demographic of FGI Participants (E1–6).
[…]
The table titled ”Demographic of FGI Participants (E1–6)” presents data about six participants, coded E1 to E6, in a focus group interview. The table is structured with five columns and seven rows, including the header row. The header row, shaded in light gray, contains the column titles: ”Code,” ”Gender,” ”Age,” ”Job Title,” and ”Years of Experience. Code Gender Age Job title Years of Experience E1 F 34 Clinical psychologist 2 years E2 F 30 Clinical psychologist 1 year E3 F 38 Clinical psychologist 9 years E4 F 36 Clinical psychologist 11 years E5 M 35 Psychiatrist 2 years E6 M 40 Psychiatrist 10 years
3.2. Findings from the Interviews
3.2.1. Challenges in Eliciting Responses from Patients with Depression
@@
Participants also noted that they had their patients engaged in paper-based diary writing methods but most demonstrated low participation rates and low engagement: “We tried a diary method on paper(in the inpatient ward), and several patients did write. What we saw was quite trivial, like, ‘I just felt bad today.’ But we learned there were significant events upon consultation, like having a big argument with other patients, which they did not record. Because patients with depression, or those who have had suicidal or self-harming incidents, often have a dulled state in expressing their emotions or feel apathetic, they tend to find such expressions very difficult.” (E3)
3.2.2. LLMs as a Bridge for Enhanced Patient Communication
Our participants envisioned LLMs as a transformative tool in mental health care, particularly for enhancing interactions with patients who struggle to express themselves. They recognized that the natural and flexible conversational abilities of LLMs could bridge communication gaps, offering a more nuanced understanding of patients’ conditions. This could be particularly beneficial in cases where patients have difficulty articulating their feelings due to symptoms like apathy or social phobia. Additionally, participants noted that using LLMs could be significantly more interactive and engaging than traditional paper-based approaches, potentially increasing compliance and participation in the therapeutic process.
Participants especially underscored the importance of capturing the continuum of thoughts leading up to a particular emotional state, such as fear, in the journaling process. They envisioned the need for using LLMs to introspect deeper into the patient’s psyche, revealing underlying thoughts and emotions that the patient might not be consciously aware of. E5 mentioned, ”It would be good if the journal continuously records the flow of thoughts. For example, it would be beneficial to document the various thoughts and detailed reasons leading up to certain feelings like fear. Like, ’I feel scared when I’m in a place with many people,’ and then digging deeper into ’Why do I feel scared?’—I think a process that gets more specific like this would be good.” This approach not only aids in a more comprehensive self-examination but also enriches the therapeutic dialogue between the patient and the MHP.
3.2.3. LLMs for Analytical Insights and Personalized Mental Health Support
The participants further suggested that LLMs could analyze journal entries to identify key themes, words, or sentiments expressed over time, offering patients tangible feedback on their emotional patterns and progress. Such analytical capabilities could empower patients with a greater sense of control and awareness of their mental health journey, potentially motivating them towards self-management and active participation in their treatment. Additionally, the analysis could assist MHPs in a deeper understanding of their patient’s emotional states and thought processes by examining the tone, choice of words, and speech or writing patterns. The participants envisioned that insights derived from LLMs about patient journaling habits could inform MHPs about the most effective counseling approaches for each individual. They suggested,  ”Observing how patients react to different forms of communication can provide valuable information. Some patients might find solace in simple reassurance, while others may benefit from more straightforward, targeted feedback.”
3.3. Improvements after the Interviews
@@
Based on the lessons from the FGI, we refined the initial concept of MindfulDiary. We leveraged the conversational abilities of LLMs to help patients document their daily experiences between clinical visits. MHPs had access to the collected data to inform their clinical decision-making. Furthermore, both MHPs and the research team concur that LLMs should not act solely as the primary intervention due to their inherent limitations but should function as supportive tools for clinical consultations. The subsequent section outlines the design and development process of our system.
@@ 4. MindfulDiary
Jane, diagnosed with chronic anxiety, frequently grapples with panic attacks. To keep track of her daily experiences, her psychiatrist recommends trying MindfulDiary as part of her treatment plan.
Every evening, Jane converses with the MindfulDiary app regarding her daily activities, emotions, and thoughts. The AI leads the conversation with Jane by asking prompted and follow-up questions about her day. After a session, the app summarizes the dialogue into a journal-style essay, on which she can revisit and reflect later. She can explore the summarized essays whenever she wants to reflect on past events or thoughts.
@@
4.1. MindfulDiary App
The MindfulDiary app for patients aims to support people who might have difficulty journaling due to apathy and cognitive load through naturalistic conversation driven by an LLM. The app consists of a home screen containing an introduction and guide to the system (Figure 2), a journal writing screen (Figure 2, 2), and a screen to review the diary entries (Figure 2).
@@ 4.1.1. Journaling User Interface
Sequential visual representation of the MindfulDiary’s journaling session flow, depicted from left to right across four cards: (1) ’Pre-Journaling Assessment’: Illustration of users taking a modified PHQ-9 mental health survey prior to their session with MindfulDiary. (2) ’User-MindfulDiary Conversation’: Image of users engaging in a conversation with MindfulDiary, narrating and documenting their day. (3) ’Summary Presentation’: Graphic depicting a diary-style summary of the three-turn conversation, with an option for users to edit. It signifies users’ choice to either continue or end the conversation. (4) ’Session Closure’: Visualization of the session’s end screen, where MindfulDiary showcases the day’s mental health assessment alongside the diary content, signifying the conclusion of the journaling experience.
Figure 3 illustrates the overall use flow of the journaling session, which begins with a Pre-Journaling Assessment (Figure 3-1⃝) that asks to fill out a questionnaire for mental health. The questionnaire comprised the modified PHQ-9 (Kroenke et al., 2001) and a custom open-ended question inquiring about recent attempts of self-harm or suicide. This assessment prevents users who provided any clues of suicidal or self-harm from journaling on the same day. (We cover this feature in detail in Section 5.4.)
@@ 4.1.2. Conversation Design
As a result, we designed the conversation of a journaling session to follow a sequence of three stages: Rapport building, Exploration, and Wrap-up. The Rapport Building state is an ice-breaker, centered on casual exchanges about a user’s day. In this state, the assistant also shares bits of information to encourage users’ openness. This approach is based on previous research findings that a chatbot’s self-disclosure positively impacts user disclosure (Lee et al., 2020a) and leverages the natural story-building ability of LLMs (Zhao et al., 2023). Overall, in this stage, our goal is to create an environment where users can comfortably share their stories. As we progress to the Exploration state, the emphasis shifts to a comprehensive understanding of the user’s daily events, feelings, and thoughts, facilitated by a mix of open-ended and closed-ended queries that ensure users remain engaged and in control of the dialogue. While open-ended queries are intended to facilitate increase the expression of feelings and emotion and less judgemental, closed-ended queries is for specific and detailed description of the experiences (Othmer and Othmer, 2002a, b).The conversation then transitions to the Wrap-up, emphasizing completion and ensuring users have fully voiced their experiences while the system remains empathetic and receptive to any lingering topics.
Besides the three main stages, we also incorporated the Sensitive Topic state that handles the most sensitive subjects, such as self-harm or suicidal ideation. When this state is triggered, psychiatrists receive instant notifications. This allows them to oversee the conversation in real-time and step in to assist the patient if necessary. Here, the system begins by empathizing with the user, recognizing their struggles, and offering a reassuring message. Following this, the system gently probes the depth of their suicidal or self-harm thoughts. If the user expresses intense or specific plans related to self-harm or suicide, the system urges them to seek prompt assistance, either at a hospital or via the local helpline.
@@ 4.1.3. Conversational Pipeline
Lengthy and complex input prompts for LLMs are known to cause poor task performance (Brown et al., 2020) by partly omitting latent concepts (Wu et al., 2022). To steer the LLM to comply with the conversational design we intended diligently, we designed MindfulDiary’s dialogue system as a state machine. Each conversation stage is carried on with a dedicated input prompt, which is more succinct and clear than a single master prompt containing instructions for all stages.
@@
Visual representation of the MindfulDiary’s conversational prompt pipeline, segmented into three distinct modules from left to right: (1) ’User Interaction’: Depicts users engaging with and responding to MindfulDiary’s messages. (2) ’Dialogue Analyzer’: Shows the intake of the recent turn count, current state, and the entirety of the user-MindfulDiary dialogue. This module processes the conversation dynamics and is built upon the GPT-4. (3) ’Response Generator’: Illustrates how the results from the Dialogue Analyzer, alongside a specified state prompt, summarization of the overall dialogue context, and the most recent three turns of the conversation, are inputted to produce a fitting response. This generated response is subsequently presented to the user. This module also operates on the GPT-4.
Figure 4 illustrates our conversation pipeline that runs each time a new user message is received. The pipeline incorporates two LLM-driven modules, a dialogue analyzer and the response generator.
The dialogue analyzer handles the stage transition, returning the stage suggestion—whether to stay or move to a new stage—and a summary paragraph of the current dialogue from the current dialogue information. The dialogue analyzer receives an input prompt that consists of the current number of turns in the conversation (progress level), the most recent stage information, and a list of criteria for recommending each stage (See 2⃝ in Figure 4). Based on this information, the underlying LLM performs a summarization task that yields a summarized paragraph of the current dialogue, a recommendation for the next stage based on the summary, turns, and the most recent stage information. For example, the system decides to move to the Wrap-up stage when the user expresses a desire to conclude or say goodbye.
The system then formulates an LLM prompt, combining a dedicated prompt for the current stage, the dialogue summary, and the recent six messages (i.e., three turn pairs) (See 3⃝ in Figure 4). Receiving the prompt as an input, the response generator generates an AI message. The stage prompt consists of the description of the task that the LLM is supposed to perform in the current stage, and the speaking rules describing the attitude that the module exhibits in the conversation. For example, the task description of Exploration stage instructs to “ask questions that encourage users to reflect on their personal stories regarding daily events, thoughts, emotions, challenges, and etc.” The speaking rules for the Rapport-building stage instruct to keep conversations simple and friendly and reply in an empathetic way.
@@ 4.3. Clinician Dashboard
Journals. This section displays the content of the journals written by patients. The information is presented in a card format, where each card offers a summary of the journal, including timestamps, total time taken to write the journal, and associated PHQ-9 score. The interaction logs between the patient and MindfulDiary are also provided in this section.
Insights. To assist professionals in browsing through the diary, this section visualizes (1) a word cloud to understand frequent terms that the participant used at a glance, (2) a summary of major events to highlight significant happenings and (3) summary of emotions to gauge the mood based on user input. When a specific period is selected for review, a comprehensive summary is generated. We used GPT-4 for most summarization tasks. To generate the word frequency data for the word cloud, we combined GPT-4 and a Korean morphological analysis package named Kiwi (Lee, 2022) to filter only nouns and verbs from the GPT output. Due to the limitations of language model-driven analysis, there might be occasional inaccuracies in the generated content. First-time users of this interface are alerted about possible inaccuracies. An in-interface tooltip also reminds users that the summarized outcomes might not be accurate.
@@ 5. Field Deployment Study
Using MindfulDiary, we conducted a four-week field deployment study with 28 patients undergoing outpatient treatment. Through the study, we aimed to explore how patients and MHPs utilize MindfulDiary and what opportunities and challenges arise from its real-world use. The study protocol was approved by the Institutional Review Board of a university hospital.
@@ 5.1. Recruitment
We targeted outpatients from the Department of Mental Health at a University Hospital. Participants were selected based on specific criteria: (1) those who had been diagnosed with MDD and (2) those who did not exhibit heightened impulsive tendencies or harbor specific intentions towards self-harm or suicide. Key exclusion criteria included a history of psychotic disorders, substance-related disorders, neurodevelopmental disorders, and neurological disorders. Eligible participants were identified through evaluations conducted by psychiatrists. Flyers and consent forms were distributed to eligible patients. For minors, the consent form process was adhered to only when they were accompanied by a guardian at the hospital.
[…]
We compensated participants on a weekly basis of participation: For participating every seven days from the starting date, participants received 15,000 KRW (approx. 11 USD). If they completed the entire four-week study process, they received 20,000 KRW as a bonus (i.e., 80,000 KRW—approx. 60 USD—in total). We did not tie the number of dialogue entries to the compensation to ensure natural data entry behavior.
As a minimum requirement for study completion, we instructed the participants not to miss four consecutive days without conversing with MindfulDiary. If a participant missed three consecutive days, an experimenter sent a reminder. In cases where participants did not respond to these reminders, their participation in the study was discontinued. This procedure was implemented to ensure active monitoring and communication. Considering that our system is designed for individuals with mental health challenges, it was crucial to maintain contact with participants and ensure their adherence to the study protocol.
Initially, 36 patients started using MindfulDiary. During the deployment, eight dropped out as they did not meet the minimum data collection requirement. These participants were disengaged from MindfulDiary due to the lack of time or decreased interest. As a result, 28 participants (P1–28; 11 males and 17 females) completed the 4-week field study and were included in the analysis. The majority of participants were adolescents and adults, with ages ranging from 12 to 28 years, with an average age of 17.6 ($SD=3.26$). Table 2 presents the demographic details and severity of depressive symptoms of the study participants. These scores are derived from psychiatric evaluations conducted within one week before the starting dates.
5.2. Procedure
Figure 5 illustrates the procedure of the field deployment study. All interviews took place remotely on Zoom.
5.2.1. MindfulDiary App
We deployed the MindfulDiary app to our patient participants. The patient protocol consisted of three parts: (1) an introductory session, (2) deployment, and (3) interviews.
Introductory Session. We first invited each participant to a remote introductory session. A researcher went through our study goal, the motivation of the MindfulDiary system, and the overall procedure of the study. We then played a demo video demonstrating how to use the MindfulDiary app. The session took about 15 minutes.
Deployment. The day following the introductory session, participants started using MindfulDiary for four weeks. We instructed participants to engage with the app whenever they have anything noteworthy but encouraged them to use it at the end of the day. We collected all data from their interactions with the MindfulDiary and the raw input content and outputs from the LLM. We asked our participants to fill out online surveys three times, at the beginning of Week 1, after Week 2, and after the deployment, to measure participants’ mental health status and their self-help capability in managing their mental health. The surveys utilized the PHQ-9 (Kroenke et al., 2001), GAD-7 (Spitzer et al., 2006), and Coping Strategies Scale (Zhao et al., 2022). (The survey results from the scale are outside the scope of this investigation.)